OpenStudy (anonymous):

2(8a + 7c - 10a + 5b) + (16a - 6(b + 15c) + 3c) in algebraic expression is 12a + 4b - 80c si,no?

razor99 (razor99):

are u asking whetere your answer is correct or no @suzy

OpenStudy (anonymous):

16a+14c-20a+10b+16a-6b-90c+3c

OpenStudy (anonymous):

im asking if its correct

razor99 (razor99):

@Nazyma everyone is not perfect okay

OpenStudy (anonymous):

@nazyma don't be so condescending

OpenStudy (anonymous):

@suzy4321 can you finish off the above line

razor99 (razor99):

srry suzy i dont think your answer is correct

OpenStudy (anonymous):

alright il go back to look at the steps again thanks though

OpenStudy (anonymous):

hint the c bit is not correct

razor99 (razor99):

i can show u all the steps suzy

razor99 (razor99):

2(8a+7c-10a+5b)+(16a-6(b+15c)+3c) -4*a+14*c+10*b+16*a-6*b-87*c -4*a+16*a+14*c-87*c+10*b-6*b 12*a-73*c+4*b

OpenStudy (anonymous):

yeah the last part i messed up anyways cant do nothing bout it i turned it in that way

OpenStudy (anonymous):

2(8a + 7c - 10a + 5b) + (16a - 6(b + 15c) + 3c)= =16a +14c-20a+10b + 16a - 6b-90c+3c= =a(16-20+16)+b(10-6)+c(14-90+3)= =12a+4b-73c

razor99 (razor99):

and suzy dont write the * sign when you do copy it okay

OpenStudy (anonymous):

*?

OpenStudy (anonymous):

o ok urs i thought u meant mine okay thanks all of u

razor99 (razor99):

no problem we are all here to help